assume that cars are equally likely to be manufactured on monday,tuesday,wednesday,thursday and friday .cars manufactured on monday have a 4% chance of being lemons .cars made on tuesday,wednesday,thursday have a 1% chance of being lemons and cars made on friday have 2% of being a lemon .if you bought a car and it turned out to be a lemon , what is the probability that was manufactured on monday

Grade:12th pass

1 Answers

kanithi premsai
6 years ago
Let everyday products be ‘y’
Then,
          Monday's lemon production be = y ×(4/100)
          Tuesday ,Wednesday ,Thursday lemon production=3 y ×(1/100)
          Friday's lemon production= y×(2/100)
          
          Total number of lemon products= y[(4/100)+(3/100)+(2/100)=y×(9/100)
 
         P(E)=NUMBER OF FAVOURABLE OUTCOMES/NUMBER OF OUTCOMES
        
         P(E)= number of lemons in Monday/no of lemons
                =[ y×(4/100)]/[y×(9/100)]
                = 4/9
 
Therefore, the probability is 4/9